what 2 numbers add to -2 but multiply to -15
Answer by etutorworld(19)

what 2 numbers add to -2 but multiply to -15
Let the numbers be x , y
As per given condition
x + y = -2............. Eq 1
x.y = -15............. Eq 2
From Eq 1 we get
y = -x -2
We plug this value of y in Eq 2
x (-x-2) = -15
-x^2 -2x = -15
Rearranging the above equation we get
x^2+ 2x - 15 = 0
Factorizing the above equation we get
(x+ 5)(x-3) = 0
Either
x +5 = 0
x = -5
Hence y = -x-2
= -(-5) -2
= 5-2
y = 3
Or x -3 = 0
x = 3
Hence y = -x-2
= -3-2
y = -5
The two numbers can be ( -5 and 3) or ( 3, -5) [ANSWER]